# Generalizing Reward Modeling for Out-of-Distribution Preference Learning

### v1

@@ Appendix C Supporting Lemmas
Proof.
@@
$=$ $\left\|\frac{1}{\beta}\exp\left(\frac{r_{\phi}(x,y)}{\beta}\right)\nabla_{\phi}r_{\phi}(x,y)\nabla_{\theta}\log\pi_{\theta^{\prime}}(y|x)-\frac{1}{\beta}\exp\left(\frac{r_{\phi}(x,y)}{\beta}\right)\nabla_{\phi}r_{\phi}(x,y)\nabla_{\theta}\log\pi_{\theta}(y|x)\right\|$
$\overset{(i)}{\leq}$ $\left\|\frac{1}{\beta}\exp\left(\frac{r_{\phi}(x,y)}{\beta}\right)\nabla_{\phi}r_{\phi}(x,y)\right\|\left\|\nabla_{\theta}\log\pi_{\theta^{\prime}}(y|x)-\nabla_{\theta}\log\pi_{\theta}(y|x)\right\|$
$\overset{(ii)}{\leq}$ $\frac{BTL}{\beta\exp\beta}\left\|\theta-\theta^{\prime}\right\|,$
where $(i)$ follows from Cauchy-Schwarz inequality and $(ii)$ follows from Assumption 2 and 3. Similarly, we also have
@@
where $(i)$ follows from Assumption 2 and 3. ∎
Appendix D Proof of the Main Theorem
D.1 Useful Lemmas
Lemma 3.
For any $k\in\{0,1,2,\ldots,K-1\}$, we have
$\begin{split}&\mathbb{E}\left[\left\|\frac{\partial\mathcal{L}_{\rm PL}({\phi_{k}},{\theta_{k,D}})}{\partial\phi_{k}}-\frac{\partial\mathcal{L}_{\rm PL}({\phi_{k}},{\theta_{*}})}{\partial\phi_{k}}\right\|^{2}\right]\\ {\leq}&8M^{2}\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{D}\frac{T^{2}}{\beta^{2}}\left(\frac{M^{2}}{L^{2}}+\left\|\theta_{0}-\theta_{*}\right\|^{2}\right)\\ &+8M^{2}\left(\frac{\alpha^{2}\sigma^{2}T^{2}}{\beta^{2}}+6\alpha^{2}\sigma_{1}^{2}+\frac{6M^{2}T^{2}}{\beta^{2}L^{2}}\sigma_{2}^{2}\right)\sum_{t=0}^{D-1}\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{t}\end{split}$ (34)
Proof.
$\overset{(i)}{\leq}$ $2M\left\|\frac{\partial{\theta_{k,D}}}{\partial\phi_{k}}-\frac{\partial{\theta_{*}}}{\partial\phi_{k}}\right\|+2L\left\|\frac{\partial{\theta_{*}}}{\partial\phi_{k}}\right\|\|\theta_{k,D}-{\theta_{*}}\|$
where $(i)$ using Assumption 2 and the fact that
$\|\mathcal{L}_{\rm PL}({\phi_{k}},{\theta_{k,D}})\|$ $=\left\|\mathbb{E}\left[\nabla\log\pi_{\theta_{k,D}}(y|x)-\nabla\log\pi_{\theta_{k,D}}(y^{\prime}|x)\right]\right\|$
$\leq\mathbb{E}\left[\left\|\nabla\log\pi_{\theta_{k,D}}(y|x)\right\|\right]+\mathbb{E}\left[\left\|\nabla\log\pi_{\theta_{k,D}}(y^{\prime}|x)\right\|\right]\leq 2M;$
$\left\|\nabla_{{\theta}}\mathcal{L}_{\rm PL}({\phi_{k}},{\theta_{k,D}})-\nabla_{{\theta}}\mathcal{L}_{\rm PL}({\phi_{k}},{\theta_{*}})\right\|$
$=$ $\left\|\mathbb{E}\left[\nabla\log\pi_{\theta_{k,D}}(y|x)-\nabla\log\pi_{\theta_{k,D}}(y^{\prime}|x)\right]-\mathbb{E}\left[\nabla\log\pi_{\theta_{*}}(y|x)-\nabla\log\pi_{\theta_{*}}(y^{\prime}|x)\right]\right\|$
$\leq$ $\mathbb{E}\left[\left\|\nabla\log\pi_{\theta_{k,D}}(y|x)-\nabla\log\pi_{\theta_{*}}(y|x)\right\|\right]+\mathbb{E}\left[\left\|\nabla\log\pi_{\theta_{k,D}}(y^{\prime}|x)-\nabla\log\pi_{\theta_{*}}(y^{\prime}|x)\right\|\right]$
$\leq$ $2L\|\theta_{k,D}-{\theta_{*}}\|,$
@@ D.1 Useful Lemmas
$\left\|\frac{\partial{\theta_{*}}}{\partial\phi_{k}}\right\|=$ $\left\|\nabla_{\phi}\nabla_{{\theta}}\mathcal{L}_{\rm FT}({\phi_{k}},{\theta_{*}})\left[\nabla_{{\theta}}^{2}\mathcal{L}_{\rm FT}({\phi_{k}},{\theta_{*}})\right]^{-1}\right\|$
@@
$\overset{(i)}{\leq}$ $\frac{1}{\beta}\mathbb{E}\left[\left\|\nabla_{{\theta}}\log\pi_{\theta_{*}}(y|x)\right\|\left\|\nabla_{\phi}r_{\phi_{k}}(x,y)\right\|\right]\mathbb{E}\left[\left\|\nabla^{2}_{{\theta}}\log\pi_{\theta_{*}}(y|x)\right\|^{-1}\right]\overset{(ii)}{\leq}\frac{MT}{\beta L},$
where $(i)$ follows from Cauchy-Schwarz’s inequality and Jensen’s inequality, $(ii)$ follows from Assumption 2. Thus, we have
$\left\|\frac{\partial\mathcal{L}_{\rm PL}({\phi_{k}},{\theta_{k,D}})}{\partial\phi_{k}}-\frac{\partial\mathcal{L}_{\rm PL}({\phi_{k}},{\theta_{*}})}{\partial\phi_{k}}\right\|\leq$ $2M\left\|\frac{\partial{\theta_{k,D}}}{\partial\phi_{k}}-\frac{\partial{\theta_{*}}}{\partial\phi_{k}}\right\|+2L\left\|\frac{\partial{\theta_{*}}}{\partial\phi_{k}}\right\|\|\theta_{k,D}-{\theta_{*}}\|$
$\leq$ $2M\left\|\frac{\partial{\theta_{k,D}}}{\partial\phi_{k}}-\frac{\partial{\theta_{*}}}{\partial\phi_{k}}\right\|+\frac{2MT}{\beta}\|\theta_{k,D}-{\theta_{*}}\|$
Taking expectation on both sides and using Lemma 4, we have Eq. (34). ∎
Lemma 4.
For any $k\in\{0,1,2,\ldots,K-1\}$,
$\begin{split}&\mathbb{E}\left[\left\|\frac{\partial\theta_{k,D}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]+\frac{T^{2}}{\beta^{2}}\mathbb{E}\left[\left\|\theta_{k,D}-\theta_{*}\right\|^{2}\right]\\ \leq&\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{D}\frac{T^{2}}{\beta^{2}}\left(\frac{M^{2}}{L^{2}}+\left\|\theta_{0}-\theta_{*}\right\|^{2}\right)\\ &+\left(\frac{\alpha^{2}\sigma^{2}T^{2}}{\beta^{2}}+6\alpha^{2}\sigma_{1}^{2}+\frac{6M^{2}T^{2}}{\beta^{2}L^{2}}\sigma_{2}^{2}\right)\sum_{t=0}^{D-1}\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{t}\end{split}$ (36)
Proof.
@@
$\mathbb{E}\left[\left\|\frac{\partial\theta_{k,t+1}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]\leq$ $3\mathbb{E}\left[\left\|I-\alpha\nabla^{2}_{{\theta}}\ell_{\rm FT}({\phi_{k}},{\theta_{k,t}};z_{k,t})\right\|^{2}\left\|\frac{\partial\theta_{k,t}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]$
@@
$\overset{(i)}{\leq}$ $3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\mathbb{E}\left[\left\|\frac{\partial\theta_{k,t}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]+\frac{6\alpha^{2}B^{2}T^{2}}{\beta^{2}\exp 2\beta}\left(L^{2}+\frac{M^{2}\rho^{2}}{L^{2}}\right)\mathbb{E}\left[\left\|\theta_{k,t}-\theta_{*}\right\|^{2}\right]$
$+6\alpha^{2}\sigma_{1}^{2}+\frac{6M^{2}T^{2}}{\beta^{2}L^{2}}\sigma_{2}^{2},$
@@
$\mathbb{E}\left[\left\|\nabla_{\phi}\nabla_{{\theta}}\ell_{\rm FT}({\phi_{k}},{\theta_{k,t}};z)-\nabla_{\phi}\nabla_{{\theta}}\mathcal{L}_{\rm FT}({\phi_{k}},{\theta_{*}})\right\|^{2}\right]\leq$ $2\mathbb{E}\left[\left\|\nabla_{\phi}\nabla_{{\theta}}\ell_{\rm FT}({\phi_{k}},{\theta_{k,t}};z)-\nabla_{\phi}\nabla_{{\theta}}\ell_{\rm FT}({\phi_{k}},{\theta_{*}};z)\right\|^{2}\right]$
@@
where $(i)$ and $(ii)$ follow from Lemma 2. Thus, we have
$\mathbb{E}\left[\left\|\frac{\partial\theta_{k,t+1}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]+\frac{2T^{2}}{\beta^{2}}\mathbb{E}\left[\left\|\theta_{k,t+1}-\theta_{*}\right\|^{2}\right]$
@@
$\leq$ $3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\mathbb{E}\left[\left\|\frac{\partial\theta_{k,t}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]+\frac{6\alpha^{2}B^{2}T^{2}}{\beta^{2}\exp 2\beta}\left(L^{2}+\frac{M^{2}\rho^{2}}{L^{2}}\right)\mathbb{E}\left[\left\|\theta_{k,t}-\theta_{*}\right\|^{2}\right]+\frac{T^{2}}{\beta^{2}}\mathbb{E}\left[\left\|\theta_{k,t+1}-\theta_{*}\right\|^{2}\right]$
$+6\alpha^{2}\sigma_{1}^{2}+\frac{6M^{2}T^{2}}{\beta^{2}L^{2}}\sigma_{2}^{2}$
$\overset{(i)}{\leq}$ $3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\mathbb{E}\left[\left\|\frac{\partial\theta_{k,t}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]+\left(\frac{6\alpha^{2}B^{2}T^{2}}{\beta^{2}\exp 2\beta}\left(L^{2}+\frac{M^{2}\rho^{2}}{L^{2}}\right)+\frac{T^{2}}{\beta^{2}}\left(1-\frac{2\alpha b\mu BL}{(b\mu+BL)\exp\beta}\right)\right)\mathbb{E}\left[\left\|\theta_{k,t}-\theta_{*}\right\|^{2}\right]$
$+\frac{\alpha^{2}\sigma^{2}T^{2}}{\beta^{2}}+6\alpha^{2}\sigma_{1}^{2}+\frac{6M^{2}T^{2}}{\beta^{2}L^{2}}\sigma_{2}^{2}$
@@
$\overset{(ii)}{\leq}$ $3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\left(\mathbb{E}\left[\left\|\frac{\partial\theta_{k,t}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]+\frac{T^{2}}{\beta^{2}}\mathbb{E}\left[\left\|\theta_{k,t}-\theta_{*}\right\|^{2}\right]\right)+\frac{\alpha^{2}\sigma^{2}T^{2}}{\beta^{2}}+6\alpha^{2}\sigma_{1}^{2}+\frac{6M^{2}T^{2}}{\beta^{2}L^{2}}\sigma_{2}^{2},$
where $(i)$ follows from Lemma 5, $(ii)$ follows from a constrain on $\beta$ such that $\beta\leq\log\frac{\alpha B^{2}}{2b\mu}\left(L^{2}+\frac{M^{2}\rho^{2}}{L^{2}}\right)$ Telescoping over $t$ from 00 to $D-1$, we have
@@
$\begin{split}&\mathbb{E}\left[\left\|\frac{\partial\theta_{k,t+1}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]+\frac{2T^{2}}{\beta^{2}}\mathbb{E}\left[\left\|\theta_{k,t+1}-\theta_{*}\right\|^{2}\right]\\ {\leq}&\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{D}\left(\left\|\frac{\partial\theta_{0,k}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}+\frac{T^{2}}{\beta^{2}}\left\|\theta_{0,k}-\theta_{*}\right\|^{2}\right)\\ &+\left(\frac{\alpha^{2}\sigma^{2}T^{2}}{\beta^{2}}+6\alpha^{2}\sigma_{1}^{2}+\frac{6M^{2}T^{2}}{\beta^{2}L^{2}}\sigma_{2}^{2}\right)\sum_{t=0}^{D-1}\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{t}\end{split}$ (39)
Applying the initialization $\theta_{k,0}=\theta_{0}$ at each inner-loop updating interation, we have the result in Eq. (36). ∎
Lemma 5.
For any $k\in\{0,1,\ldots,K-1\}$ and any $t\in\{0,1,\ldots,D-1\}$,
Proof.
Taking expectation on both sides conditioning on $\theta_{k,t}$, we have

### v2

@@ Appendix 0.C Supporting Lemmas
$\left\|\nabla_{\phi}\nabla_{\theta}\ell_{\rm FT}(\phi,\theta)-\nabla_{\phi}\nabla_{\theta}\ell_{\rm FT}(\phi^{\prime},\theta)\right\|$ $\leq\frac{MB(P+T^{2})}{\beta\exp\beta}\|\phi-\phi^{\prime}\|$ (30)
$\left\|\nabla_{\theta}^{2}\ell_{\rm FT}(\phi,\theta)-\nabla_{\theta}^{2}\ell_{\rm FT}(\phi^{\prime},\theta)\right\|$ $\leq\frac{BTL}{\beta\exp\beta}\|\phi-\phi^{\prime}\|$ (31)
Proof
@@
$=$ $\left\|\frac{1}{\beta}\exp\left(\frac{r_{\phi}(x,y)}{\beta}\right)\nabla_{\phi}^{\top}r_{\phi}(x,y)\nabla_{\theta}\log\pi_{\theta^{\prime}}(y|x)-\frac{1}{\beta}\exp\left(\frac{r_{\phi}(x,y)}{\beta}\right)\nabla_{\phi}^{\top}r_{\phi}(x,y)\nabla_{\theta}\log\pi_{\theta}(y|x)\right\|$
[…]
$\overset{(i)}{\leq}$ $\frac{BTL}{\beta\exp\beta}\left\|\theta-\theta^{\prime}\right\|,$
where $(i)$ follows from Assumption 2 and 3.
Similarly, we also have
@@
where $(i)$ follows from Assumption 2 and 3.
Similarly, we have
$\left\|\nabla_{\phi}\nabla_{\theta}\ell_{\rm FT}(\phi,\theta)-\nabla_{\phi}\nabla_{\theta}\ell_{\rm FT}(\phi^{\prime},\theta)\right\|$
$=$ $\left\|\frac{1}{\beta}\exp\left(\frac{r_{\phi^{\prime}}(x,y)}{\beta}\right)\nabla_{\phi}^{\top}r_{\phi^{\prime}}(x,y)\nabla_{\theta}\log\pi_{\theta}(y|x)-\frac{1}{\beta}\exp\left(\frac{r_{\phi}(x,y)}{\beta}\right)\nabla_{\phi}^{\top}r_{\phi}(x,y)\nabla_{\theta}\log\pi_{\theta}(y|x)\right\|$
$\leq$ $\frac{1}{\beta}\left\|\nabla_{\theta}\log\pi_{\theta}(y|x)\right\|\left\|\exp\left(\frac{r_{\phi^{\prime}}(x,y)}{\beta}\right)\nabla_{\phi}^{\top}r_{\phi^{\prime}}(x,y)-\exp\left(\frac{r_{\phi}(x,y)}{\beta}\right)\nabla_{\phi}^{\top}r_{\phi}(x,y)\right\|$
$\leq$ $\frac{MB}{\beta\exp\beta}\left\|\nabla_{\phi}^{\top}r_{\phi^{\prime}}(x,y)-\nabla_{\phi}^{\top}r_{\phi}(x,y)\right\|+\frac{MT}{\beta}\left\|\exp\left(\frac{r_{\phi^{\prime}}(x,y)}{\beta}\right)-\exp\left(\frac{r_{\phi}(x,y)}{\beta}\right)\right\|$
$\leq$ $\frac{MB(P+T^{2})}{\beta\exp\beta}\|\phi-\phi^{\prime}\|$
Similarly, we have
$\left\|\nabla_{\theta}^{2}\ell_{\rm FT}(\phi,\theta)-\nabla_{\theta}^{2}\ell_{\rm FT}(\phi^{\prime},\theta)\right\|$
$=$ $\left\|\nabla_{\theta}^{2}\log\pi_{\theta}(y|x)\exp\left(\frac{r_{\phi^{\prime}}(x,y)}{\beta}\right)-\nabla_{\theta}^{2}\log\pi_{\theta}(y|x)\exp\left(\frac{r_{\phi}(x,y)}{\beta}\right)\right\|$
$\leq$ $\|\nabla_{\theta}^{2}\log\pi_{\theta}(y|x)\|\left\|\exp\left(\frac{r_{\phi^{\prime}}(x,y)}{\beta}\right)-\exp\left(\frac{r_{\phi}(x,y)}{\beta}\right)\right\|$
$\leq$ $\frac{BTL}{\beta\exp\beta}\|\phi-\phi^{\prime}\|$
Lemma 3
Based on Assumption 2, we have the Lipschitz property of the preference learning loss $\ell_{\rm PL}$, $\nabla_{\theta}\ell_{\rm PL}(\phi,\theta)$,
$\|\ell_{\rm PL}(\phi,\theta)-\ell_{\rm PL}(\phi,\theta^{\prime})\|\leq$ $2M\|\theta-\theta^{\prime}\|$ (32)
$\left\|\nabla_{\theta}\ell_{\rm PL}(\phi,\theta)-\nabla_{\theta}\ell_{\rm PL}(\phi,\theta^{\prime})\right\|\leq$ $(2L+M^{2})\|\theta-\theta^{\prime}\|$ (33)
Proof
For Eq. (33), we have
$\|\ell_{\rm PL}(\phi,\theta)-\ell_{\rm PL}(\phi,\theta^{\prime})\|$
$=$ $\|\log\sigma\left(\log\pi_{\theta}(y|x)-\log\pi_{\theta}(y^{\prime}|x)\right)-\log\sigma\left(\log\pi_{\theta^{\prime}}(y|x)-\log\pi_{\theta^{\prime}}(y^{\prime}|x)\right)\|$
$\overset{(i)}{\leq}$ $\|\log\pi_{\theta}(y|x)-\log\pi_{\theta}(y^{\prime}|x)-\log\pi_{\theta^{\prime}}(y|x)+\log\pi_{\theta^{\prime}}(y^{\prime}|x)\|$
$\overset{(ii)}{\leq}$ $\|\log\pi_{\theta}(y|x)-\log\pi_{\theta^{\prime}}(y|x)\|+\|\log\pi_{\theta^{\prime}}(y^{\prime}|x)-\log\pi_{\theta}(y^{\prime}|x)\|$
$\overset{(iii)}{\leq}$ $2M\|\theta-\theta^{\prime}\|,$
@@
where $(i)$ follows from the $1$-Lipschitz of $\log\sigma(\cdot)$, $(ii)$ follows from the triangle inequality and $(iii)$ follows from Assumption 2.
For Eq. (3), based on the triangle inequality, we have
$\left\|\nabla_{\theta}\ell_{\rm PL}(\phi,\theta)-\nabla_{\theta}\ell_{\rm PL}(\phi,\theta^{\prime})\right\|$
$=$ $\|\sigma(\log\pi_{\theta}(y^{\prime}|x)-\log\pi_{\theta}(y|x))\left(\nabla_{\theta}\log\pi_{\theta}(y|x)-\nabla_{\theta}\log\pi_{\theta}(y^{\prime}|x)\right)$
$\quad-\sigma(\log\pi_{\theta^{\prime}}(y^{\prime}|x)-\log\pi_{\theta^{\prime}}(y|x))\left(\nabla_{\theta}\log\pi_{\theta^{\prime}}(y|x)-\nabla_{\theta}\log\pi_{\theta^{\prime}}(y^{\prime}|x)\right)\|$
$\leq$ $\|\sigma(\log\pi_{\theta}(y^{\prime}|x)-\log\pi_{\theta}(y|x))\left(\nabla_{\theta}\log\pi_{\theta}(y|x)-\nabla_{\theta}\log\pi_{\theta}(y^{\prime}|x)-\nabla_{\theta}\log\pi_{\theta^{\prime}}(y|x)+\nabla_{\theta}\log\pi_{\theta^{\prime}}(y^{\prime}|x)\right)\|$
$+\|\left(\sigma(\log\pi_{\theta}(y^{\prime}|x)-\log\pi_{\theta}(y|x))-\sigma(\log\pi_{\theta^{\prime}}(y^{\prime}|x)-\log\pi_{\theta^{\prime}}(y|x))\right)\left(\nabla_{\theta}\log\pi_{\theta^{\prime}}(y|x)-\nabla_{\theta}\log\pi_{\theta^{\prime}}(y^{\prime}|x)\right)\|$
$\overset{(i)}{\leq}$ $\|\nabla_{\theta}\log\pi_{\theta}(y|x)-\nabla_{\theta}\log\pi_{\theta^{\prime}}(y|x)\|+\|\nabla_{\theta}\log\pi_{\theta}(y^{\prime}|x)-\nabla_{\theta}\log\pi_{\theta^{\prime}}(y^{\prime}|x)\|$
$+\frac{M}{2}\left(\|\log\pi_{\theta}(y^{\prime}|x)-\log\pi_{\theta^{\prime}}(y^{\prime}|x)\|+\|\log\pi_{\theta}(y|x)-\log\pi_{\theta^{\prime}}(y|x)\|\right)$
$\overset{(ii)}{\leq}$ $(2L+M^{2})\|\theta-\theta^{\prime}\|,$
where the first term of $(i)$ follows from the fact w.r.t. logistic funtion that $\sigma(\cdot)<1$, the second term of $(i)$ follows from the $\frac{1}{4}$-lipschitz of logistic funtion $\sigma$ and Assumption 2, $(ii)$ follows from Assumption 2.
Lemma 4
Based on Assumption 2, for any $k\in\{0,1,2,\ldots,K-1\}$, we have
$\left\|\frac{\partial{\theta_{*}}}{\partial\phi_{k}}\right\|\leq\frac{MT}{\beta\mu}$ (34)
Proof
@@
$\left\|\frac{\partial{\theta_{*}}}{\partial\phi_{k}}\right\|$
$=$ $\left\|\nabla_{\phi}\nabla_{{\theta}}\mathcal{L}_{\rm FT}({\phi_{k}},{\theta_{*}})\left[\nabla_{{\theta}}^{2}\mathcal{L}_{\rm FT}({\phi_{k}},{\theta_{*}})\right]^{-1}\right\|$
@@
$\overset{(i)}{\leq}$ $\frac{1}{\beta}\mathbb{E}\left[\left\|\nabla_{\phi}^{\top}r_{\phi_{k}}(x,y)\right\|\left\|\nabla_{{\theta}}\log\pi_{\theta_{*}}(y|x)\right\|\right]\mathbb{E}\left[\left\|\nabla^{2}_{{\theta}}\log\pi_{\theta_{*}}(y|x)\right\|^{-1}\right]\overset{(ii)}{\leq}\frac{MT}{\beta\mu},$
where $(i)$ follows from Jensen’s inequality, $(ii)$ follows from Assumption 1, 2.
Appendix 0.D Proof of the Main Theorem
0.D.1 Useful Lemmas
Lemma 5
Based on Assumption 1, 2 and 3, $\ell_{\rm PL}(\phi,\theta)$ ($\mathcal{L}_{\rm PL}(\phi,\theta)$) is $P_{\phi}$-smoothness w.r.t. $\phi$, where
$P_{\phi}=\frac{4M^{3}B(P+T^{2})}{\beta b\mu}+\frac{(6L+M^{2})B^{2}T^{2}M^{2}}{\beta^{2}b^{2}\mu^{2}}+\frac{2\rho B^{3}T^{2}M^{3}}{\beta^{2}b^{3}\mu^{3}}$ (36)
Proof
The smoothness of $\ell_{\rm PL}(\phi,\theta)$ is adopted from Lemma 2.2 in [17] with Lemma 1, 2 and 3. Then, the smoothness of $\mathcal{L}_{\rm PL}(\phi,\theta)$ follows from Jensen’s inequality.
Lemma 6
Based on Assumption 1, 2 and 3, we have for any $k\in\{0,1,\ldots,K-1\}$ and any $t\in\{0,1,\ldots,D-1\}$,
Proof
Taking expectation on both sides conditioning on $\theta_{k,t}$, we have
$\mathbb{E}\left[\left\|\theta_{k,t+1}-\theta_{*}\right\|^{2}\right]$
$\overset{(i)}{\leq}$ $\alpha^{2}\left(\sigma^{2}+\left\|\nabla_{\theta}\mathcal{L}_{\rm FT}\left({\phi_{k}},\theta_{k,t}\right)\right\|^{2}\right)-2\alpha\langle\nabla_{\theta}\mathcal{L}_{\rm FT}\left({\phi_{k}},\theta_{k,t}\right),\theta_{k,t}-\theta_{*}\rangle+\left\|\theta_{k,t}-\theta_{*}\right\|^{2}$
where $(i)$ follows from the definition of the variance ${\rm Var[X]}=\mathbb{E}[X^{2}]-\mathbb{E}[X]^{2}$, $(ii)$ follows from the fact that the $\frac{b\mu}{\exp\beta}$-strong convexity and $\frac{BL}{\exp\beta}$-smoothness of $\mathcal{L}_{\rm FT}({\phi},{\theta})$ w.r.t. $\theta$ in Lemma 1 deduces that $\langle\nabla_{\theta}\mathcal{L}_{\rm FT}\left({\phi_{k}},\theta_{k,t}\right)-\nabla_{\theta}\mathcal{L}_{\rm FT}\left({\phi_{k}},\theta_{*}\right),\theta_{k,t}-\theta_{*}\rangle\geq\frac{b\mu BL}{(b\mu+BL)\exp\beta}\|\theta_{k,t}-\theta_{*}\|^{2}+\frac{\exp\beta}{b\mu+BL}\|\nabla_{\theta}\mathcal{L}_{\rm FT}\left({\phi_{k}},\theta_{k,t}\right)\|^{2}$ (Theorem 2.1.12 in [25]), where $\nabla_{\theta}\mathcal{L}_{\rm FT}\left({\phi_{k}},\theta_{*}\right)=0$.
Using the condition $\alpha\leq\frac{2\exp\beta}{b\mu+BL}$ and unconditioning on $\theta_{k,t-1}$, we have Eq. (37).
Lemma 7
Based on Assumption 1, 2 and 3, we have for any $k\in\{0,1,2,\ldots,K-1\}$,
$\begin{split}&\mathbb{E}\left[\left\|\frac{\partial\theta_{k,D}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]+\frac{T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}\mathbb{E}\left[\left\|\theta_{k,D}-\theta_{*}\right\|^{2}\right]\\ \leq&\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{D}\frac{T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}\left(\frac{4M^{2}}{(2L+M^{2})^{2}}+\left\|\theta_{0}-\theta_{*}\right\|^{2}\right)\\ &+\left(\frac{\alpha^{2}\sigma^{2}T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}+6\alpha^{2}\left(\sigma_{1}^{2}+\frac{M^{2}T^{2}}{\beta^{2}\mu^{2}}\sigma_{2}^{2}\right)\right)\sum_{t=0}^{D-1}\left(3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\right)^{t}\end{split}$ (38)
Proof
@@ 0.D.1 Useful Lemmas
$\mathbb{E}\left[\left\|\frac{\partial\theta_{k,t+1}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]$
$\leq$ $3\mathbb{E}\left[\left\|I-\alpha\nabla^{2}_{{\theta}}\ell_{\rm FT}({\phi_{k}},{\theta_{k,t}};z_{k,t})\right\|^{2}\left\|\frac{\partial\theta_{k,t}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]$
@@
$\overset{(i)}{\leq}$ $3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\mathbb{E}\left[\left\|\frac{\partial\theta_{k,t}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]+\frac{6\alpha^{2}B^{2}T^{2}}{\beta^{2}\exp 2\beta}\left(L^{2}+\frac{M^{2}\rho^{2}}{\mu^{2}}\right)\mathbb{E}\left[\left\|\theta_{k,t}-\theta_{*}\right\|^{2}\right]$
$+6\alpha^{2}\left(\sigma_{1}^{2}+\frac{M^{2}T^{2}}{\beta^{2}\mu^{2}}\sigma_{2}^{2}\right),$
@@
$\mathbb{E}\left[\left\|\nabla_{\phi}\nabla_{{\theta}}\ell_{\rm FT}({\phi_{k}},{\theta_{k,t}};z)-\nabla_{\phi}\nabla_{{\theta}}\mathcal{L}_{\rm FT}({\phi_{k}},{\theta_{*}})\right\|^{2}\right]$
$\leq$ $2\mathbb{E}\left[\left\|\nabla_{\phi}\nabla_{{\theta}}\ell_{\rm FT}({\phi_{k}},{\theta_{k,t}};z)-\nabla_{\phi}\nabla_{{\theta}}\ell_{\rm FT}({\phi_{k}},{\theta_{*}};z)\right\|^{2}\right]$
@@
where $(i)$ and $(ii)$ follow from Lemma 2.
Thus, we have
$\mathbb{E}\left[\left\|\frac{\partial\theta_{k,t+1}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]+\frac{T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}\mathbb{E}\left[\left\|\theta_{k,t+1}-\theta_{*}\right\|^{2}\right]$
@@
$\leq$ $3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\mathbb{E}\left[\left\|\frac{\partial\theta_{k,t}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]+\frac{6\alpha^{2}B^{2}T^{2}}{\beta^{2}\exp 2\beta}\left(L^{2}+\frac{M^{2}\rho^{2}}{\mu^{2}}\right)\mathbb{E}\left[\left\|\theta_{k,t}-\theta_{*}\right\|^{2}\right]$
@@
$+\frac{T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}\mathbb{E}\left[\left\|\theta_{k,t+1}-\theta_{*}\right\|^{2}\right]+6\alpha^{2}\left(\sigma_{1}^{2}+\frac{M^{2}T^{2}}{\beta^{2}\mu^{2}}\sigma_{2}^{2}\right)$
@@
$\overset{(i)}{\leq}$ $3\left(1-\frac{\alpha b\mu}{\exp\beta}\right)^{2}\mathbb{E}\left[\left\|\frac{\partial\theta_{k,t}}{\partial\phi_{k}}-\frac{\partial\theta_{*}}{\partial\phi_{k}}\right\|^{2}\right]+\frac{\alpha^{2}\sigma^{2}T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}+6\alpha^{2}\left(\sigma_{1}^{2}+\frac{M^{2}T^{2}}{\beta^{2}\mu^{2}}\sigma_{2}^{2}\right)$
$+\left(\frac{6\alpha^{2}B^{2}T^{2}}{\beta^{2}\exp 2\beta}\left(L^{2}+\frac{M^{2}\rho^{2}}{\mu^{2}}\right)+\frac{T^{2}(2L+M^{2})^{2}}{4\beta^{2}\mu^{2}}\left(1-\frac{2\alpha b\mu BL}{(b\mu+BL)\exp\beta}\right)\right)\mathbb{E}\left[\left\|\theta_{k,t}-\theta_{*}\right\|^{2}\right]$